OpenStudy (issy14):

Derive: w = Square root of (x^2 · 5^x )3 : So this is what I did: ((x^2 + 5^x)^3)^1/2 = (x^2+5^x) = 3/2(x^2 + 5^x)^1/2 (2x* 5^x + ln(5))

OpenStudy (issy14):

BOB ( Bac of the book) has a different answer, here is my question, do I have to take the derivative again of (x^2 + 5^x) ?

OpenStudy (issy14):

back*

OpenStudy (issy14):

Is it because it's a product rule ? I think that's what I was missing.

OpenStudy (anonymous):

hey :)

OpenStudy (anonymous):

wait is that a dot or+ between the two terms ??

OpenStudy (anonymous):

@Issy14

OpenStudy (issy14):

it's a , which two terms?

OpenStudy (issy14):

it's a * times (multiplication)

OpenStudy (anonymous):

so if it was a multiplication why did u write it as addition ?

OpenStudy (issy14):

\[w=\sqrt{(x^{2}*5^x)^3}\]

OpenStudy (issy14):

that's what it is

OpenStudy (issy14):

that's the original.

OpenStudy (anonymous):

\[\large \sqrt{(x^2 · 5^x )^3}\] Is this your question?

OpenStudy (issy14):

sorry, wrong sign all the way

OpenStudy (anonymous):

ok first u have to write it like this \[(x ^{2}.5^{x})^{\frac{ 3 }{ 2}}\]

OpenStudy (anonymous):

Is there \(\cdot\) or \(+\) ??

OpenStudy (issy14):

so this is what i came up so far with: 3/2(x^2* 5^x) (2x)(5^x) + (5^x * ln(5) (x^2)

OpenStudy (anonymous):

yup u wrote it as addition in the second step then u derived it as addition and that's why ur answer is wrong

OpenStudy (issy14):

I truly apologize for the mishap

OpenStudy (anonymous):

You have to subtract also in the power, 3/2 you are carrying forward, and then you will do (3/2) - 1 in the power..

OpenStudy (issy14):

I did that, at the beginning: I ended up with 1/2 so 3(x^2*5^x)^1/2 and I proceeded from there

OpenStudy (anonymous):

Wait, please write your final answer or the step where you have reached till now..

OpenStudy (issy14):

3/2(x^2* 5^x) (2x)(5^x) + (5^x * ln(5) (x^2)

OpenStudy (anonymous):

Why you are not taking exponent there???

OpenStudy (issy14):

exponent there in 3/2(x^2* 5^x)^1/2 (2x)(5^x) + (5^x * ln(5) (x^2)

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

And please be careful with brackets.. At last, you have applied product rule which contains two terms, so first term upto ..........^1/2, after this brackets will start...

OpenStudy (issy14):

got it, but then the back of the book has a different answer and I don't know how they got there

OpenStudy (anonymous):

\[\large \color{green}{\implies \frac{3}{2}(x^2* 5^x)^{\frac{1}{2}} [(2x)(5^x) + (5^x * \ln(5) (x^2)]}\]

OpenStudy (anonymous):

Sorry, in excitement, I wrote somewhat big there...!!! :)

OpenStudy (issy14):

Yes, that's what I have so far, that's where I stopped. I don't know how to proceed from there.

OpenStudy (issy14):

hahahahhaha it's ok, I like math too, it makes think and be patient.

OpenStudy (anonymous):

Now, tell me what your BOB, I mean what your back of book says.. Please write that answer too..

OpenStudy (tkhunny):

Just for the record, are you SURE that "derive" is a imperative form synonymous with "find the derivative"?

OpenStudy (issy14):

ok, one second BOB says: 3/2 x^2 square root of 5^2x (2 + x ln(5))

OpenStudy (issy14):

If they factored, idk where they got the common factor from

OpenStudy (anonymous):

Is that 5^(2x) there??

OpenStudy (issy14):

yes =(

OpenStudy (anonymous):

And you should be happy that you have done absolutely right...

OpenStudy (issy14):

so it's BOB crazy?

OpenStudy (issy14):

nuuuuu, don't say that, sometimes BOB has been known to be wrong

OpenStudy (anonymous):

Oh sorry, you must be happy now that we are absolutely and perfectly right... Sorry, I did not read the BOB answer carefully... Yes we are right, just we have to rearrange our reached answer..

OpenStudy (issy14):

hahhahahaha...

OpenStudy (anonymous):

See, this you know: \[\large (a \cdot b)^x = a^x \cdot b^x\]

OpenStudy (anonymous):

And now just take \(x \cdot 5^x\) common from the brackets... And see what you get..

OpenStudy (anonymous):

And you must be knowing: \[\large (a)^{\frac{1}{2}} = \sqrt{a}\]

OpenStudy (issy14):

5x? shouldn't it be 5^x that's what you meant right?

OpenStudy (anonymous):

\[\large (x^2 \cdot 5^x)^{\frac{1}{2}} = (x^2)^{\frac{1}{2}} \cdot (5^x)^{\frac{1}{2}}\]

OpenStudy (issy14):

ok, one second let's see.

OpenStudy (anonymous):

Where???

OpenStudy (anonymous):

Ha ha ha ha... :P

OpenStudy (anonymous):

Where you have seen 5x ??? Open your eyes, there is no 5x anywhere... :)

OpenStudy (issy14):

omg, I think I finally lost it.

OpenStudy (anonymous):

See do step by step..

OpenStudy (issy14):

BOB was right once again... it's a hate-love relationship. 3/2 x^2 square root 5^x ( 2 + ln(5))

OpenStudy (anonymous):

You got it or not??

OpenStudy (issy14):

Yeah I got it. I wrote it on my whiteboard

OpenStudy (anonymous):

Okay, that's nice.. :) Good Luck...!!!

OpenStudy (anonymous):

And Well Done!!!

OpenStudy (issy14):

Thank you!!! my brain is sore like a muscle, I'm going to drink water.

OpenStudy (anonymous):

Okay!!

OpenStudy (tkhunny):

Except that \(\sqrt{x^{2}} = |x|\) in the absence of additional information.